# Self-Dual Codes and LCD Codes in Sum-Rank Metric

### v1

@@ 4.1 The First Method
In this subsection, we introduce the first method based on Subsection 3.1. Using this method, we can construct self-dual sum-rank codes and LCD sum-rank codes in $\mathbb{F}_{2}^{(2,2),\ldots,(2,2)}$. The following is an important theorem from Lemma 3.1.
@@
Since $C_{0}$, $C_{1}$ are linear codes over $\mathbb{F}_{4}$ of length $t$, according to Theorem 2.15, $SR(C_{0},C_{1})$ is a linear sum-rank code with block length $t$ and matrix size $2\times 2$. By Lemma 3.1, we have,
@@
Using this theorem, we can construct some self-dual sum-rank codes or LCD sum-rank codes over $\mathbb{F}_{2}$ with block length $t$ and matrix size $2\times 2$ from two Euclidean self-dual codes or Euclidean LCD codes over $\mathbb{F}_{4}$ with length $t$. In addition, according to Theorem 2.15, to some extent, the larger the minimum Hamming distances of $C_{0}$ and $C_{1}$ are, the larger the minimum sum-rank distance of $SR(C_{0},C_{1})$ will be.
When $C_{0}=C_{1}$, we can determine $d_{sr}(SR(C_{0},C_{1}))$.
@@
Let $C\subseteq\mathbb{F}_{4}^{t}$ be a $[t,k,d]_{4}$ code. Then the minimum sum-rank distance of $SR(C,C)$
@@
Since the minimum Hamming distance of $C$ is $d$, there exist a codeword $\bm{c}=(c_{0},\ldots,c_{t-1})\in C$ such that $d_{H}(\bm{c})=d$. Let $w$ be a primitive element of $\mathbb{F}_{4}$. Then $w^{2}=w+1$ and $\{1,w\}$ is an ordered basis of $\mathbb{F}_{4}$ over $\mathbb{F}_{2}$. Let $c_{i}=c_{i}^{(1)}+c_{i}^{(2)}w$, where $c_{i}^{(1)}$, $c_{i}^{(2)}\in\mathbb{F}_{2}$, $10\leq i\leq t-$ 1. Then the linear map $c_{i}x+c_{i}x^{2}$ over $\mathbb{F}_{4}$ corresponds to the matrix
@@
If $c_{i}=0$, $M_{i}=\bm{0}$, giving $rank(M_{i})=0$. If $c_{i}\neq 0$, either $c_{i}^{(1)}$ or $c_{i}^{(2)}$ is not equal to 00, giving $rank(M_{i})=1$. Therefore, $wt_{sr}(SR(\bm{c},\bm{c}))=d$. Hence, $d_{sr}(SR(C,C))=d.$
@@
Reference [28] provides a database of the best known Euclidean self-dual codes over $\mathbb{F}_{4}$ (with largest minimum Hamming distances). We can construct many self-dual sum-rank codes with good parameters from these codes. For example, there is only one $[2,1,2]$ self-dual code over $\mathbb{F}_{4}$, so we let both $C_{0}$ and $C_{1}$ be the code. Then by Theorem 2.15 and Theorem 4.4, $SR(C_{0},C_{1})$ is a self-dual sum-rank code with block length $2$, matrix size $2\times 2$ and dimension $4$. By the proposition above, we find that its minimum sum-rank distance $d_{sr}(SR(C_{0},C_{1}))=2$. We present those with block length up to $30$ in Table 1. In Table 1, we list the length of $C_{0}$ and $C_{1}$ (block length of $SR(C_{0},C_{1})$) in column 1 and the minimum Hamming distance of $C_{0}$ and $C_{1}$ in column 2. In the last two columns, we present the dimension as well as the minimum sum-rank distance of $SR(C_{0},C_{1})$. We note that it is not necessary to let $C_{0}=C_{1}$, because in some cases there exist more than one best known Euclidean self-dual codes and when $C_{0}\neq C_{1}$, $SR(C_{0},C_{1})$ has larger minimum sum-rank distance. (Even if $C_{0}$ is monomial equivalent to $C_{1}$, we think that they are two different codes.)
@@
In addition, we can construct cyclic self-dual sum-rank codes of block length larger than $30$. By Lemma 2.6, Theorem 2.15, Theorem 4.4 and Proposition 4.6, we can get the theorem below.
@@
Let $t=\frac{2(2^{2m}-1)}{u}$, where $m=3,5,7,\ldots$ and $u$ is a divisor of $12^{2m}-$ 1. Then there exists a cyclic self-dual sum-rank code over $\mathbb{F}_{2}$ with block length $t$, matrix size $2\times 2$, dimension $2t$ and minimum sum-rank distance at least $\sqrt{\frac{2t}{u}}-\frac{4}{u}$.
@@
[…]
@@
[…]
@@
As for the construction of LCD sum-rank codes, it seems to be easier than self-dual sum-rank codes. According to Theorem 2.3, only if there exists an $[n,k,d]_{4}$ linear code, can we find an $[n,k,d]_{4}$ Euclidean LCD code. Then we can use those Euclidean LCD codes with good parameters to get some LCD sum-rank codes. For example, we can get the following family of distance-optimal LCD sum-rank codes.
@@
Let $l\geq 3$. Then there exist two Euclidean LCD codes over $\mathbb{F}_{4}$, $C_{0}$, $C_{1}$ with parameters $[4^{l}-1,4^{l}-2,2]$ and $[4^{l}-1,4^{l}-2l-2,4]$ respectively. Furthermore, $SR(C_{0},C_{1})\subseteq\mathbb{F}_{2}^{(2,2),\ldots,(2,2)}$ is a distance-optimal LCD sum-rank code with block length $14^{l}-$ 1, dimension $4(4^{l}-l-2)$ and minimum sum-rank distance $4$.
Combining Theorem 2.4, Lemma 2.16 and Theorem 4.4, we can prove the theorem directly.
Similarly, we can construct some cyclic LCD sum-rank codes from those classical cyclic LCD codes. Since there are too many classical cyclic LCD codes of various length constructed by different authors, we cannot list them one by one. Thus, we only give some examples. The following two examples are from some related results in [23] and the computation by MAGMA.
@@
At last, we present some cyclic LCD sum-rank codes of larger block length. According to Lemma 2.7, it is easy to get the following lemma.
@@
    $C_{(4,\frac{4^{m}+1}{5},2\cdot 4^{\frac{m-1}{2}}+1,1)}$ has parameters $[\frac{4^{m}+1}{5},\frac{4^{m}+1}{5}-12m(4^{\frac{m-3}{2}}-1),\geq 2\cdot 4^{\frac{m-1}{2}}+1]$.
@@
    $C_{(4,\frac{4^{m}+1}{5},2(\cdot 4^{\frac{m-1}{2}}+1),0)}$ has parameters $[\frac{4^{m}+1}{5},\frac{4^{m}+1}{5}-12m(4^{\frac{m-3}{2}}-1)-1,\geq 4^{\frac{m+1}{2}}+2]$.
@@
    $C_{(4,\frac{4^{m}+1}{5},3\cdot 4^{\frac{m-1}{2}}+1,1)}$ has parameters $[\frac{4^{m}+1}{5},\frac{4^{m}+1}{5}-2m(9\cdot 4^{\frac{m-3}{2}}-14),\geq 3\cdot 4^{\frac{m-1}{2}}+1]$.
@@
    $C_{(4,\frac{4^{m}+1}{5},3\cdot 4^{\frac{m-1}{2}}+2,0)}$ has parameters $[\frac{4^{m}+1}{5},\frac{4^{m}+1}{5}-2m(9\cdot 4^{\frac{m-3}{2}}-14)-1,\geq 6\cdot 4^{\frac{m-1}{2}}+2]$.
@@ 4.2 The Second Method
By Theorem 2.4, if $q$ is even or both $q$ and $m$ are odd, a self-dual basis $\mathcal{B}$ of $\mathbb{F}_{q^{m}}$ over $\mathbb{F}_{q}$ exists. Due to Corollary 3.7, we can get other results,
@@
According to Lemma 3.4, $d_{sr}(Mat_{\mathcal{B}}(C))=d_{sr}(C)$. However, $d_{sr}(C)$ is difficult to get exactly. In most cases, we only know the exact value of $d_{H}(C)$. In the following, we will present upper and lower bounds of $d_{sr}(Mat_{\mathcal{B}}(C))$ (or $d_{sr}(C)$) in terms of $d_{H}(C)$.
@@
$c_{k}=\begin{cases}1,&\mbox{if }1\leq k\leq n_{s}+l;\\ 0,&\mbox{if }n_{s}+l+1\leq k\leq N.\end{cases}$
@@
Now, let $m=n_{1}=\cdots=n_{t}=2$, we can get the corollary easily.
@@
By Theorem 4.16, we connect self-dual sum-rank codes and LCD sum-rank codes with Euclidean self-dual codes and Euclidean LCD codes. Using this connection, we can construct some self-dual sum-rank codes and LCD sum-rank codes in $\mathbb{F}_{q}^{(m,n_{1}),\ldots,(m,n_{t})}$ provided $q$ is even or both $q$ and $m$ are odd. What is more, according to Proposition 4.17 and Corollary 4.18, the larger the minimum Hamming distance of $C$ is, the larger the minimum sum-rank distance of $Mat_{\mathcal{B}}(C)$ will be.
@@
Let $q=m=n_{1}=\cdots=n_{t}=2$. We can obtain some self-dual sum-rank codes with those Euclidean self-dual codes over $\mathbb{F}_{4}$ with the largest minimum Hamming distances in [28]. For example, let $C$ be the unique $[2,1,2]$ Euclidean self-dual code over $\mathbb{F}_{4}$ and =$\mathcal{B}=\{w,w^{2}\}$ be the self-dual basis of $\mathbb{F}_{4}$ over $\mathbb{F}_{2}$ where $w$ is a primitive element of $\mathbb{F}_{4}$. (Throughout the subsection, we keep the notation.) Then, by Theorem 4.16, $Mat_{\mathcal{B}}(C)$ is a self-dual sum-rank code with block length $1$, matrix size $2\times 2$ and dimension $2$. By computation, its minimum sum-rank distance is equal to $1$. We present those with block length up to $15$ in Table 7. In this table, we list the block length, the dimension and the minimum sum-rank distance of $Mat_{\mathcal{B}}(C)$ in columns 1, 3 and 4. In column 2, we give the minimum Hamming distance of $C$.
@@
$1$ $2$ $2$ $d_{sr}(Mat_{\mathcal{B}}(C))=1$
$2$ $3$ $4$ $d_{sr}(Mat_{\mathcal{B}}(C))=2$
$3$ $3$ $6$ $2\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 3$
$4$ $4$ $8$ $2\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 4$
$5$ $4$ $10$ $2\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 4$
$6$ $6$ $12$ $3\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 6$
$7$ $6$ $14$ $3\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 6$
$8$ $6$ $16$ $3\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 6$
$9$ $6$ $18$ $3\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 6$
$10$ $8$ $20$ $4\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 8$
$11$ $8$ $22$ $4\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 8$
$12$ $8$ $24$ $4\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 8$
$13$ $8$ $26$ $4\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 8$
$14$ $9$ $28$ $5\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 9$
@@
It is obvious that $Mat_{\mathcal{B}}(C)$ is cyclic if $C$ is cyclic, so we can also construct some cyclic self-dual sum-rank codes with those known cyclic self-dual codes over $\mathbb{F}_{4}$ just like we did in the first method.
@@
Then we consider those cyclic self-dual sum-rank codes with larger length. Using Lemma 2.6, Lemma 3.4, Theorem 4.16 and Corollary 4.18, we derive the following result.
@@
Let $t=\frac{2^{2m}-1}{u}$, where $m=3,5,7,\ldots$ and $u$ is a divisor of $12^{2m}-$ 1. Then there exists a family of cyclic self-dual sum-rank codes over $\mathbb{F}_{2}$ with block length $t$, matrix size $2\times 2$, dimension $2t$ and minimum sum-rank distance at least $\sqrt{\frac{t}{u}}-\frac{2}{u}$.
@@
Table 8: Cyclic LCD Sum-Rank Codes $Mat_{\mathcal{B}}(C)$ of Block Length $6$
@@
Let $m>5$ be an odd integer and $Mat_{\mathcal{B}}\,:\,\mathbb{F}_{4}^{\frac{4^{m}+1}{5}}\rightarrow\mathbb{F}_{2}^{(2,3),(2,2),\ldots,(2,2)}$ be the $\mathbb{F}_{2}$-linear isomorphism between $\mathbb{F}_{4}^{\frac{4^{m}+1}{5}}$ and $\mathbb{F}_{2}^{(2,3),(2,2),\ldots,(2,2)}$ where =$\mathcal{B}=\{w,w^{2}\}$ is the self-dual basis of $\mathbb{F}_{4}$ over $\mathbb{F}_{2}$. We have the following $4$ cyclic LCD sum-rank codes in $\mathbb{F}_{2}^{(2,3),(2,2),\ldots,(2,2)}$ with block length $\frac{2(4^{m-1}-2)}{5}$:

### v2

@@ 4.1 The First Method
In this subsection, we introduce the first method based on Subsection 3.1. Using this method, we can construct self-dual sum-rank codes and LCD sum-rank codes in $\mathbb{F}_{2}^{(2,2),\ldots,(2,2)}$. The following is the key theorem from Theorem 3.4.
@@
Since $C_{0}$, $C_{1}$ are linear codes over $\mathbb{F}_{4}$ of length $t$, according to Theorem 3.3, $SR(C_{0},C_{1})$ is a linear sum-rank code with block length $t$ and matrix size $2\times 2$. By Theorem 3.4, we have
@@
Using this theorem, we can construct some self-dual sum-rank codes or LCD sum-rank codes over $\mathbb{F}_{2}$ with block length $t$ and matrix size $2\times 2$ from two Euclidean self-dual codes or Euclidean LCD codes in $\mathbb{F}_{4}^{t}$ . In addition, according to Theorem 3.3, to some extent, the larger the minimum Hamming distances of $C_{0}$ and $C_{1}$ are, the larger the minimum sum-rank distance of $SR(C_{0},C_{1})$ will be.
When $C_{0}=C_{1}$, we can determine the exact value of $d_{sr}(SR(C_{0},C_{1}))$.
@@
Let $C\subseteq\mathbb{F}_{4}^{t}$ be a $t,k,d]_{4}$ code. Then the minimum sum-rank distance of $SR(C,C)$
@@
Since the minimum Hamming distance of $C$ is $d$, there exists a codeword $\bm{c}=(c_{0},\ldots,c_{t-1})\in C$ such that $d_{H}(\bm{c})=d$. Let $w$ be a primitive element of $\mathbb{F}_{4}$. Then $w^{2}=w+1$ and $\{1,w\}$ is an ordered basis of $\mathbb{F}_{4}$ over $\mathbb{F}_{2}$. Let $c_{i}=c_{i}^{(1)}+c_{i}^{(2)}w$, where $c_{i}^{(1)}$, $c_{i}^{(2)}\in\mathbb{F}_{2}$, $0\leq i\leq t-1$. Then the linear map $c_{i}x+c_{i}x^{2}$ over $\mathbb{F}_{4}$ corresponds to the matrix
@@
If $c_{i}=0$, $M_{i}=\bm{0}$, giving $rank(M_{i})=0$. If $c_{i}\neq 0$, either $c_{i}^{(1)}$ or $c_{i}^{(2)}$ is not equal to 0, giving $rank(M_{i})=1$. Thus, $wt_{sr}(SR(\bm{c},\bm{c}))=d$. Hence, $d_{sr}(SR(C,C))=d.$
@@
Reference [19] provides a table of the best known Euclidean self-dual codes over $\mathbb{F}_{4}$ (with largest minimum Hamming distances). We can construct many self-dual sum-rank codes with good parameters from these codes. For example, there is only one $[2,1,2]$ self-dual code over $\mathbb{F}_{4}$, so we let both $C_{0}$ and $C_{1}$ be the code. Then by Theorem 3.3 and Theorem 4.4, $SR(C_{0},C_{1})$ is a self-dual sum-rank code with block length $2$, matrix size $2\times 2$ and dimension $4$. By the proposition above, we find that its minimum sum-rank distance $d_{sr}(SR(C_{0},C_{1}))=2$. We present those with block length up to $30$ in Table 1. We list the length of $C_{0}$ and $C_{1}$ (the block length of $SR(C_{0},C_{1})$) in column 1 and the minimum Hamming distance of $C_{0}$ and $C_{1}$ in column 2. In the last two columns, we present the dimension as well as the minimum sum-rank distance of $SR(C_{0},C_{1})$. We note that it is not necessary to let $C_{0}=C_{1}$, because in some cases there exist more than one best known Euclidean self-dual codes and when $C_{0}\neq C_{1}$, $SR(C_{0},C_{1})$ has larger minimum sum-rank distance. (Even if $C_{0}$ is monomial equivalent to $C_{1}$, we think that they are two different codes.)
@@
In addition, we can construct cyclic self-dual sum-rank codes of block length larger than $30$. To achieve this goal, we need some cyclic self-dual codes of larger and more flexible lengths. The following is an example.
Lemma 4.9.
([11, Theorem 4.1]) Let $n=\frac{2(2^{sm}-1)}{u}$, where $m=3,5,7,\ldots$ and $u$ is a divisor of $2^{sm}-1$. Then there exists a family of Euclidean self-dual repeated-root cyclic codes over $\mathbb{F}_{2^{s}}$ with the length $n$ and the minimum distance at least $\sqrt{\frac{2^{s-1}n}{u}}-\frac{2^{s}}{u}$.
By Theorem 3.3, Theorem 4.4, Proposition 4.6 and Lemma 4.9, we can get the cyclic self-dual code in sum-rank metric below.
@@
Let $t=\frac{2(2^{2m}-1)}{u}$, where $m=3,5,7,\ldots$ and $u$ is a divisor of $2^{2m}-1$. Then there exists a cyclic self-dual sum-rank code over $\mathbb{F}_{2}$ with block length $t$, matrix size $2\times 2$, dimension $2t$ and minimum sum-rank distance at least $\sqrt{\frac{2t}{u}}-\frac{4}{u}$.
@@
Let $C_{i}\subseteq\mathbb{F}_{4}^{t}$ be a $t,k_{i},d_{i}]_{4}$ self-dual code, i=0, 1i=0,\,1. Then the minimum sum-rank distance of $SR(C_{0},C_{1})$ is at most
@@
Since $C_{i}$ is a $t,k_{i},d_{i}]_{4}$ self-dual code, i=0, 1i=0,\,1, by Lemma 2.2, $d_{i}\leq 4\lfloor\frac{t}{12}\rfloor+4$. Thus,
@@
As for the construction of LCD sum-rank codes, it seems to be easier than the construction of self-dual sum-rank codes. According to Theorem 2.3, only if there exists an $n,k,d]_{4}$ linear code, can we find an $n,k,d]_{4}$ Euclidean LCD code. Then we can use those Euclidean LCD codes with good parameters to get some LCD sum-rank codes.
If $C\subseteq\mathbb{F}_{q}^{(m_{1},n_{1}),\ldots,(m_{t},n_{t})}$ is a sum-rank code with $M$ codewords and the minimum sum-rank distance $d_{sr}$ and there is no sum-rank code in $\mathbb{F}_{q}^{(m_{1},n_{1}),\ldots,(m_{t},n_{t})}$ with $M$ codewords and the minimum sum-rank distance $d_{sr}+1$, we call $C$ a distance-optimal sum-rank code. The authors of [12] constructed a family of distance-optimal cyclic sum-rank codes in $\mathbb{F}_{2}^{(2,2),\ldots,(2,2)}$ as below.
Lemma 4.12.
([12, Corollary 7.3]) Let $l\geq 3$, $C_{0}$ be the cyclic code of length $4^{l}-1$ over $\mathbb{F}_{4}$ with generator polynomial x1x-1 and $C_{1}$ be the BCH code $C_{(4,4^{l}-1,4,0)}$. Then $SR(C_{0},C_{1})\subseteq\mathbb{F}_{2}^{(2,2),\ldots,(2,2)}$ has dimension $4(4^{l}-l-2)$ and minimum sum-rank distance $4$. Furthermore, the sum-rank cyclic code $SR(C_{0},C_{1})$ is distance-optimal.
Correspondingly, we can get the following family of distance-optimal LCD sum-rank codes.
@@
Let $l\geq 3$. Then there exist two Euclidean LCD codes over $\mathbb{F}_{4}$, $C_{0}$, $C_{1}$ with parameters $4^{l}-1,4^{l}-2,2]$ and $4^{l}-1,4^{l}-2l-2,4]$ respectively. Furthermore, $SR(C_{0},C_{1})\subseteq\mathbb{F}_{2}^{(2,2),\ldots,(2,2)}$ is a distance-optimal LCD sum-rank code with block length $4^{l}-1$, dimension $4(4^{l}-l-2)$ and minimum sum-rank distance $4$.
Combining Theorem 2.3, Theorem 4.4 and Lemma 4.12, we can prove the theorem directly.
Apart from this, we can construct some cyclic LCD sum-rank codes from those classical cyclic LCD codes. We know that a cyclic code of length $n$ over $\mathbb{F}_{q}$ is an LCD code if 1-1 is a power of $q$ modulo $n$ (see [28]), so cyclic codes over $\mathbb{F}_{q}$ with length $n=\frac{q^{m}+1}{N}$ are always LCD codes, where $1\leq N\leq q^{m}+1$ is a divisor of $q^{m}+1$. There are many cyclic LCD codes over $\mathbb{F}_{q}$ with length $n=\frac{q^{m}+1}{N}$. In [15], the authors provided some relevant examples. The following two examples are from some related results in [15] and the computation by MAGMA.
@@
At last, we construct some cyclic LCD sum-rank codes of larger block lengths with the lemma below.
Lemma 4.16.
([15, Theorem 3.4]) Suppose $n=\frac{q^{m}+1}{q+1}$, where $m>5$ is odd. For $\delta=lq^{\frac{m-1}{2}}+1$ with $2\leq l\leq q-1$, the code $C_{(q,n,\delta,1)}$ and $C_{(q,n,\delta+1,0)}$ have parameters $n,k,\geq\delta]$ and $n,k-1,\geq 2\delta]$, respectively, where
Using the lemma above, we can get the following four cyclic LCD codes over $\mathbb{F}_{4}$.
@@
    $C_{(4,\frac{4^{m}+1}{5},2\cdot 4^{\frac{m-1}{2}}+1,1)}$ has parameters $\frac{4^{m}+1}{5},\frac{4^{m}+1}{5}-12m(4^{\frac{m-3}{2}}-1),\geq 2\cdot 4^{\frac{m-1}{2}}+1]$.
@@
    $C_{(4,\frac{4^{m}+1}{5},2(\cdot 4^{\frac{m-1}{2}}+1),0)}$ has parameters $\frac{4^{m}+1}{5},\frac{4^{m}+1}{5}-12m(4^{\frac{m-3}{2}}-1)-1,\geq 4^{\frac{m+1}{2}}+2]$.
@@
    $C_{(4,\frac{4^{m}+1}{5},3\cdot 4^{\frac{m-1}{2}}+1,1)}$ has parameters $\frac{4^{m}+1}{5},\frac{4^{m}+1}{5}-2m(9\cdot 4^{\frac{m-3}{2}}-14),\geq 3\cdot 4^{\frac{m-1}{2}}+1]$.
@@
    $C_{(4,\frac{4^{m}+1}{5},3\cdot 4^{\frac{m-1}{2}}+2,0)}$ has parameters $\frac{4^{m}+1}{5},\frac{4^{m}+1}{5}-2m(9\cdot 4^{\frac{m-3}{2}}-14)-1,\geq 6\cdot 4^{\frac{m-1}{2}}+2]$.
@@ 4.2 The Second Method
If $q$ is even or both $q$ and $m$ are odd, a self-dual basis $\mathcal{B}$ of $\mathbb{F}_{q^{m}}$ over $\mathbb{F}_{q}$ exists. Due to Corollary 3.9, we can get these results,
@@
According to Lemma 3.6, $d_{sr}(Mat_{\mathcal{B}}(C))=d_{sr}(C)$. However, $d_{sr}(C)$ is difficult to get exactly. In most cases, we only know the exact value of $d_{H}(C)$. Hence, we derive upper and lower bounds of $d_{sr}(Mat_{\mathcal{B}}(C))$ (or $d_{sr}(C)$) in terms of $d_{H}(C)$.
@@
$c_{k}=\begin{cases}1,&\mbox{if }1\leq k\leq n_{s}+l,\\ 0,&\mbox{if }n_{s}+l+1\leq k\leq N.\end{cases}$
@@
Now, let $m=n_{1}=\cdots=n_{t}=2$, it is easy to get the corollary as follows.
@@
By Theorem 4.19, we connect self-dual sum-rank codes and LCD sum-rank codes with Euclidean self-dual codes and Euclidean LCD codes. Using this connection, we can construct some self-dual sum-rank codes and LCD sum-rank codes in $\mathbb{F}_{q}^{(m,n_{1}),\ldots,(m,n_{t})}$ provided $q$ is even or both $q$ and $m$ are odd. What is more, according to Proposition 4.20 and Corollary 4.21, to some extent, the larger the minimum Hamming distance of $C$ is, the larger the minimum sum-rank distance of $Mat_{\mathcal{B}}(C)$ will be.
@@
Let $q=m=n_{1}=\cdots=n_{t}=2$. We can obtain some self-dual sum-rank codes with those Euclidean self-dual codes over $\mathbb{F}_{4}$ with the largest minimum Hamming distances in [19]. For example, let $C$ be the unique $[2,1,2]$ Euclidean self-dual code over $\mathbb{F}_{4}$ and $\mathcal{B}=\{w,w^{2}\}$ be the self-dual basis of $\mathbb{F}_{4}$ over $\mathbb{F}_{2}$ where $w$ is a primitive element of $\mathbb{F}_{4}$. (Throughout the subsection, we keep the notation.) Then, by Theorem 4.19, $Mat_{\mathcal{B}}(C)$ is a self-dual sum-rank code with block length $1$, matrix size $2\times 2$ and dimension $2$. By computation, its minimum sum-rank distance is equal to $1$. We present those with block length up to $15$ in Table 7. In this table, we list the block length, the dimension and the minimum sum-rank distance of $Mat_{\mathcal{B}}(C)$ in columns 1, 3 and 4. In column 2, we give the minimum Hamming distance of $C$.
@@
11 22 22 $d_{sr}(Mat_{\mathcal{B}}(C))=1$
22 33 44 $d_{sr}(Mat_{\mathcal{B}}(C))=2$
33 33 66 $2\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 3$
44 44 88 $2\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 4$
55 44 $10$ $2\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 4$
66 66 $12$ $3\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 6$
77 66 $14$ $3\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 6$
88 66 $16$ $3\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 6$
99 66 $18$ $3\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 6$
$10$ 88 2020 $4\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 8$
$11$ 88 $22$ $4\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 8$
$12$ 88 $24$ $4\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 8$
$13$ 88 $26$ $4\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 8$
$14$ 99 $28$ $5\leq d_{sr}(Mat_{\mathcal{B}}(C))\leq 9$
@@
It is evident that $Mat_{\mathcal{B}}(C)$ is cyclic if $C$ is cyclic, so we can also construct some cyclic self-dual sum-rank codes with those known cyclic self-dual codes over $\mathbb{F}_{4}$ just like we did in the first method.
@@
Then we consider those cyclic self-dual sum-rank codes with larger lengths. Using Lemma 3.6, Lemma 4.9, Theorem 4.19 and Corollary 4.21, we derive the following result.
@@
Let $t=\frac{2^{2m}-1}{u}$, where $m=3,5,7,\ldots$ and $u$ is a divisor of $2^{2m}-1$. Then there exists a family of cyclic self-dual sum-rank codes over $\mathbb{F}_{2}$ with block length $t$, matrix size $2\times 2$, dimension $2t$ and minimum sum-rank distance at least $\sqrt{\frac{t}{u}}-\frac{2}{u}$.
Remark 4.26.
Compared this theorem with Theorem 4.10, it seems that the parameters of the two classes of cyclic self-dual sum-rank codes are very similar, but in fact, the block lengths of them are definitely different. In details, the block lengths of the cyclic self-dual sum-rank codes in Theorem 4.10 are even while those of the cyclic self-dual sum-rank codes in this theorem are odd. It is shown that we can construct cyclic self-dual sum-rank codes with various lengths by using the two methods.
@@
Table 8: Cyclic LCD Sum-Rank Codes $Mat_{\mathcal{B}}(C)$ of Block Length 66
@@
Let $m>5$ be an odd integer and $Mat_{\mathcal{B}}\,:\,\mathbb{F}_{4}^{\frac{4^{m}+1}{5}}\rightarrow\mathbb{F}_{2}^{(2,3),(2,2),\ldots,(2,2)}$ be the $\mathbb{F}_{2}$-linear isomorphism between $\mathbb{F}_{4}^{\frac{4^{m}+1}{5}}$ and $\mathbb{F}_{2}^{(2,3),(2,2),\ldots,(2,2)}$ where $\mathcal{B}=\{w,w^{2}\}$ is the self-dual basis of $\mathbb{F}_{4}$ over $\mathbb{F}_{2}$. We have the following $4$ cyclic LCD sum-rank codes in $\mathbb{F}_{2}^{(2,3),(2,2),\ldots,(2,2)}$ with block length $\frac{2(4^{m-1}-2)}{5}$: